OpenStudy (anonymous):

What is the slope of the line 5x - 2y = 10?

OpenStudy (anonymous):

Slope of line is general form: -x/y -5/-2= 5/2

OpenStudy (anonymous):

Or you can: -2y= -5x +10 y=5/2x-5

OpenStudy (anonymous):

okay that was much eaier . thank You